OpenStudy (anonymous):

prove : sin A/(1+cos A) = tan 1/2 A

OpenStudy (anonymous):

multiply top and bottom by (1-cos A)

OpenStudy (anonymous):

[sin A ( 1 - cos A)] / [( 1 + cos A) ( 1 - cos A) ]

OpenStudy (anonymous):

the numerator is sin A - sin A cos A , the denominator is 1 - cos ^2 A , , which is equal to sin^2 A

OpenStudy (anonymous):

hmm, its easier if you start from the right side

OpenStudy (anonymous):

haha so are we gonna start from the right side?i have absolutely no idea how to solve.

OpenStudy (valpey):

Let b = 1/2 A \[\sin2b/(1+\cos2b) = 2\sin(b)\cos(b)/(1+\cos^2(b)-\sin^2(b))\] \[1 = \sin^2() + \cos^2()\] \[\sin2b/(1+\cos2b) = 2\sin(b)\cos(b)/(\cos^2(b)+sin^2(b)+\cos^2(b)-\sin^2(b))\] \[=2\sin(b)\cos(b)/2\cos^2(b) = \sin(b)/\cos(b) = \tan(b) = \tan(1/2A)\]

OpenStudy (anonymous):

there is a tan half formula ?

OpenStudy (anonymous):

tan (1/2 A) = sin (1/2 A ) / cos (1/2 A)

OpenStudy (lalaly):

\[sinA=2\sin(\frac{A}{2})\cos(\frac{A}{2})\]\[cosA=2\cos^2(\frac{A}{2})-1\] therfore \[\frac{2\sin(\frac{A}{2})\cos(\frac{A}{2})}{2\cos^2(\frac{A}{2})} \]\[=\frac{\sin(\frac{A}{2})}{\cos(\frac{A}{2})} = \tan(\frac{A}{2})\]

OpenStudy (anonymous):

THANKS lalaly and others for the answers.much appreciated people.

OpenStudy (lalaly):

ur welcome:)

OpenStudy (anonymous):

wait theres a mistak

OpenStudy (anonymous):

sin (a/2) = ?

OpenStudy (anonymous):

oh i see , you multiplied top and bottom by 2 cos (a/2)

OpenStudy (lalaly):

me?

OpenStudy (lalaly):

i didnt multiply anything, i used the trigonometric identities

OpenStudy (anonymous):

im coming from the right side of the equation

OpenStudy (anonymous):

that gives me the clue to use half angle formula

OpenStudy (anonymous):

but youre way works